OpenStudy (anonymous):

x-3 over x+4= x-4 over x+1....solve for x?

OpenStudy (anonymous):

(x-3)/(x+4) = (x-4)/(x+1) ===> x^2-16 =x^2-2x-3 ===> 2x=13==>x=13/2